OpenStudy (anonymous):

write as the log of a single argument: 3log(x)-1/2log(y)-4log(z)

OpenStudy (aum):

Use the following identities to simplify: \(m\log(A) = \log(A)^m\) \(\log(A) - \log(B) = \log(A/B)\)